# On compatibility of Koszul- and higher preprojective gradings

### v2

@@ Abstract.
We investigate compatibility of gradings for an almost Koszul or Koszul algebra $R$ that is also the higher preprojective algebra $\Pi_{n+1}(A)$ of an $n$-hereditary algebra $A$. For an $n$-representation finite algebra $A$, we show that $A$ must be Koszul if $\Pi_{n+1}(A)$ can be endowed with an almost Koszul grading. For a basic $n$-representation infinite algebra $A$, we show that $A$ must be Koszul if $\Pi_{n+1}(A)$ can be endowed with a Koszul grading. From this we deduce that a higher preprojective grading of an (almost) Koszul algebra $R=\Pi_{n+1}(A)$ is, in both cases, isomorphic to a cut of the (almost) Koszul grading. Up to a further assumption on the tops of the degree 00 subalgebras for the different gradings, we also show a similar result without the basic assumption in the $n$-representation infinite case. As an application, we show that $n$-APR tilting preserves the property of being Koszul for $n$-representation infinite algebras.
@@ 1. Introduction
In this way, $\Pi(Q)$ inherits a grading from tensor degrees, and one recovers $kQ$ as the degree 00 part of this grading. Furthermore, a change of orientation of $Q$ amounts to a change of the grading on $\Pi(Q)$, so while $\Pi(Q)$ as an ungraded algebra does not retain this information, the additional structure does. It is this perspective on the preprojective algebra that was generalised by Iyama and Oppermann [IO13] in Iyama’s higher Auslander-Reiten theory. Here, for a finite dimensional algebra $A$ of global dimension $n$, the $(n+1)$-preprojective algebra is defined as
@@
It is this interplay between the two gradings that we want to investigate in this article, both for the $n$-representation finite and the $n$-representation infinite case. The motivating question, which is made precise in 2.24, asks whether a given higher preprojective grading and a Koszul grading on some algebra $R$ need to be compatible in the way described above. We begin by explaining that this is not literally true and give some easy examples. However, the examples suggest that for two “incompatible” gradings we can find an automorphism that moves the higher preprojective grading to one that is compatible with the given Koszul grading. We give a proof of this statement in the $n$-representation finite case in 4.3. In the $n$-representation infinite case, we prove an analogous statement for basic algebras in 5.14, and a non-basic version over a fixed semisimple base ring in 5.15. The proof is based on the fact that if $A$ is ring-indecomposable and $n$-hereditary, and $\Pi_{n+1}(A)$ can be endowed with some (almost) Koszul grading, then $A$ can be endowed with a Koszul grading. In this way, our results can be seen as a converse to Grant-Iyama’s work, showing that $A$ can be given a Koszul grading if and only if $\Pi_{n+1}(A)$ can be given an (almost) Koszul grading. Additionally, this answers a question raised in [DG24a, Question 3.13] in the basic, $n$-representation tame case, which also applies to [DG24b]. In the $n$-representation tame case, our results have a clear geometric interpretation. The compatibility of gradings becomes equivalent to the commutation of two $k^{\ast}$-actions on the affine variety $\operatorname{Spec}(Z(\Pi_{n+1}(A)))$, as outlined in Remark 2.25. Our results also complement the work in [Thi20], where Thibault shows that certain Koszul algebras can not arise as higher preprojective algebras. Similarly, this shows that in order to construct preprojective gradings on Koszul quiver algebras, one can exhaust all gradings up to isomorphism by grading the quiver, as has been done in [HIO14, DG24a, DG24b, Gio19]. As an application, we show that $n$-APR tilting preserves Koszulity.
@@ 2.2. Graded algebras
For a graded ring $R_{\bullet}$ and graded modules $M$, $N$, we denote by
@@
the graded Hom- and Ext-spaces. That means, $\operatorname{Hom}_{R_{\bullet}}(M,N)$ consists of homogeneous morphisms of degree 00.
@@ 2.2.1. Koszul algebras
[MV07, Theorem 2.8] Let $R=\bigoplus_{i\geq 0}R_{i}$ be a locally finite dimensional, nonnegatively graded algebra with $R_{0}$ semisimple. Then $R$ is Koszul if and only if $\operatorname{Ext}^{*}_{R}(R_{0},R_{0})=\bigoplus_{i\geq 0}\operatorname{Ext}^{i}_{R}(R_{0},R_{0})$ is generated in cohomological degrees 00 and $1$.
@@
Let $A$ be a finite dimensional algebra, and let $S=A/\operatorname{rad}(A)$. Then $A$ admits a Koszul grading if and only if $\operatorname{Ext}^{*}_{A}(S,S)$ is generated in cohomological degrees 00 and $1$.
@@
Consider the polynomial ring $R=k[x]=k[x+1]$, with two gradings placing $x$ respectively $x+1$ in degree $1$. Then the degree 00 parts are not isomorphic as simple $R$-modules. Of course, the automorphism exchanging $x$ and $x+1$ is a graded isomorphism for the two gradings, so we can adjust one grading so that the degree 00 parts coincide.
@@ 2.4. The Grant-Iyama construction
[GI20, Proposition 4.16] If $A_{\bullet}$ is Koszul, then the $A$-bimodule $\operatorname{Ext}^{n}_{A}(D(A),A)$ is generated in degree $-n$ with respect to the Koszul grading.
@@ 2.5. Higher preprojective cuts
    The (almost) Koszul generating bimodule $R_{1}$ decomposes into preprojectively homogeneous pieces of degrees 00 and $1$, i.e. we have $R_{1}={}_{0}(R_{1})\oplus{}_{1}(R_{1})$ such that (${}_{i}(R_{1})\leq{}_{i}R$ as vector spaces.
@@
    The decomposition recovers the preprojective grading, i.e. we have ${}_{0}R=\langle R_{0}\oplus{}_{0}(R_{1})\rangle$ as algebras and ${}_{1}R$ is generated by (${}_{1}(R_{1})$ as an ${}_{0}R$-bimodule.
@@
We first show that each (almost) Koszul homogeneous piece $R_{i}$ decomposes into a direct sum of preprojectively homogeneous pieces. By assumption, the decomposition is true for $R_{0}$ and $R_{1}$. Furthermore, it is obvious that $R_{i}\supseteq\left(\bigoplus_{j\geq 0}R_{i}\cap{}_{j}R\right)$. Then note that $R_{i}$ is generated over $R_{0}$ by products of length $i$ of elements in $R_{1}$, and use distributivity to see that $R_{i}\subseteq\left(\bigoplus_{j\geq 0}R_{i}\cap{}_{j}R\right)$. Now, we show that a preprojectively homogeneous piece ${}_{i}R$ decomposes into (almost) Koszul homogeneous components. The argument is the same, we note ${}_{0}R=\langle R_{0}\oplus{}_{0}(R_{1})\rangle$ is generated by (almost) Koszul homogeneous elements, and the same holds for ${}_{1}R=({}_{0}R)({}_{1}(R_{1}))({}_{0}R)$. Then, writing each element in ${}_{j}R$ as a product of length $j$ of elements in ${}_{1}R$ and using distributivity shows the claim. ∎
@@
Denote by $E=\operatorname{Ext}^{n}_{A}(D(A),A)$ the bimodule which generates the preprojective algebra. Then it follows from 2.18 that $E$ as a graded $A$-module is generated in degree $-n$. One sees easily that the Koszul degree 00 piece is $R_{0}=A_{0}$, and that $R_{1}=A_{1}\oplus E_{-n}$. Since $A$ is Koszul, it follows that $A$ is generated in degree $1$, and $E$ is generated over $A$ by $R_{1}\cap E=E_{-n}$. ∎
@@
In particular, the Koszul degree $1$ piece $R_{1}=\langle a,b,a^{\ast},b^{\ast}\rangle$ differs from $\langle a,b,(a^{\ast})^{r},(b^{\ast})^{r}\rangle$, but declaring $\{a,b\}$ to have preprojective degree $1$ still defines a preprojective grading. Denote its degree 00 part by $B=\langle e+a,f-a,(a^{\ast})^{r},(b^{\ast})^{r}\rangle$. Then even the semisimple base rings differ, since $R_{0}\cap B=0$. However, we note that the graded radical for the Koszul grading is $\operatorname{grad}(R_{\bullet})=(a,b,a^{\ast},b^{\ast})$. This equals the graded radical of the newly defined higher preprojective grading, which is given by the ideal $(a,b,(a^{\ast})^{r},(b^{\ast})^{r})$. The equality is not a coincidence, as 5.6 shows.
@@
It was proven in [HIO14] that a ring-indecomposable $n$-hereditary algebra is either $n$-representation finite or $n$-representation infinite. This is reflected in the dimension of the higher preprojective algebra being finite or infinite. We therefore split the above question into the two cases of finite and infinite as well. We will give a positive answer to this question for the $n$-representation finite case. We also prove this for basic algebras in the $n$-representation infinite case. The proof hinges on the fact that the preprojective degree 00 part can be given a Koszul grading, which may differ from the (almost) Koszul grading on the preprojective algebra.
@@ 3. Reduction to the graded basic case
$eRe\simeq\operatorname{End}_{R}(Re)\text{ and }R\simeq\operatorname{End}_{eRe}(eR).$
Note that $e$ is homogeneous of degree 00. Therefore, $eRe$ is graded, and $eR$ is a graded $eRe$-module. This way, $\operatorname{End}_{eRe}(eR)$ becomes graded and $R_{\bullet}\simeq\operatorname{End}_{eRe}(eR)$ is a graded isomorphism. Next, we decompose the $eRe$-module $eR$ further as $eR=\bigoplus_{i,j}eRe_{i,j}$ according to the decomposition of $1$ in $R_{0}$. From the isotypic decomposition, we obtain that there exist elements $E_{i,j_{1},j_{2}}\in R_{0}$ such that $e_{i,j_{1}}E_{i,j_{1},j_{2}}=e_{i,j_{2}}$. Since $E_{i,j_{1},j_{2}}\in R_{0}$, we conclude that the $eRe$-modules $eRe_{i,j_{1}}$ and $eRe_{i,j_{2}}$ are graded isomorphic. Hence, we can collect the graded modules $\bigoplus_{j=1}^{d_{i}}eRe_{i,j}\simeq(eRe_{i,1})^{d_{i}}$ and obtain the graded isomorphism
@@
Let $e\in R_{0}$ be the Morita idempotent that defines the graded basic algebra $R_{\bullet}^{b}=eRe$. Decompose $e=e_{1}+\ldots+e_{m}$ inside of $R_{0}$ into primitive idempotents. It then follows from 3.6 that we have a graded isomorphism $R\simeq\operatorname{End}_{eRe}(\bigoplus_{i=1}^{m}(eRe_{i})^{d_{i}})$ for some multiplicities $d_{i}$. Since $eRe$ is graded basic, we have that $eR_{0}e$ is basic and finite dimensional. The second component of the bigrading (,)($\bullet$,$\ast$)( ∙ , ∗ ) on $eRe$ defines a grading on $eR_{0}e$. However, it follows from 3.3 that this grading is trivial on $eR_{0}e$, and hence we have $R^{b}_{(0,0)}=\bigoplus_{i\geq 0}R^{b}_{(0,i)}$. Recall that the identity in $eRe$ is $e$, which consequently has bidegree $(0,0)$, as do all the $e_{i}$. It follows that each projective $R^{b}_{\bullet}$ module $eRe_{i}$ is bigraded. Hence, $R\simeq\operatorname{End}_{eRe}(\bigoplus_{i=1}^{m}(eRe_{i})^{d_{i}})$ becomes $\mathbb{Z}^{2}$-graded as well. Finally, it follows from 3.6 that the first component of the bigrading on $\operatorname{End}_{eRe}(\bigoplus(eRe_{i})^{d_{i}})$ recovers the $\bullet$-grading on $R$. ∎
@@
The fact that the second component of the grading $R_{\bullet,\ast}$ is a higher preprojective grading follows from [Thi20, Proposition 4.12], whose proof applies for both the $n$-representation finite and the $n$-representation infinite case. To see that $R_{\ast}$ is a cut for the Koszul grading $R_{\bullet}$, we first note that clearly $R_{1}=\bigoplus_{j\geq 0}R_{(1,j)}$ decomposes into preprojectively homogeneous pieces. Furthermore, recall that any idempotent $e\in R_{0}$ affording the Morita equivalence with $R_{\bullet}^{b}$ is full. Therefore, it follows that $R_{(1,j)}=0$ for all $j\geq 2$, because the same is true for the cut on $R^{b}_{\bullet}$. To check the second property of a cut, note that $R_{0}$ becomes a graded algebra with respect to the $\ast$-component of the bigrading. By 3.3, it follows that $R_{0}=R_{(0,0)}$, so we have that $R_{0}\subseteq\bigoplus_{i\geq 0}R_{(i,0)}$ has preprojective degree 00. Since $R_{\bullet}$ is Koszul, it is generated in degrees 00 and $1$, and hence we have that $\bigoplus_{i\geq 0}R_{(i,0)}$ is generated by $R_{(0,0)}\oplus R_{(1,0)}=R_{0}\oplus R_{(1,0)}$. ∎
@@
Suppose $R_{\bullet}$ and ${}_{\bullet}R$ are a Koszul and a higher preprojective grading respectively, so that $A={}_{0}R$ is $n$-representation infinite. We may choose $e\in R_{0}$ for the reduction to the graded basic algebra $R_{\bullet}^{b}$. However, in general $e\not\in{}_{0}R$ need not be preprojectively homogeneous. This means that the subspaces $e({}_{i}R)e$ need not define a grading on $eRe$, so we can not transfer the higher preprojective grading this way. In order to remedy this, we set $f={}_{0}e$ and in this way obtain a full, homogeneous idempotent in ${}_{0}R$, which gives rise to a graded basic algebra $fRf=\bigoplus_{i\geq 0}(f({}_{i}R)f)$. However, since $R$ is infinite dimensional, the reductions $eRe$ and $fRf$ need not be isomorphic. We therefore will assume that $R_{0}\subseteq{}_{0}R$ in 5.15, so that the idempotent $e\in R_{0}$ is also homogeneous with respect to the preprojective grading. As pointed out in Example 2.23, this assumption is not fulfilled in general. In the spirit of the Wedderburn-Malcev theorem, one can hope that $R_{0}$ can be moved into ${}_{0}R$ by some automorphism, so we record the following question.

### v3

@@ Abstract.
We investigate compatibility of gradings for an almost Koszul or Koszul algebra $R$ that is also the higher preprojective algebra $\Pi_{n+1}(A)$ of an $n$-hereditary algebra $A$. For an $n$-representation finite algebra $A$, we show that $A$ must be Koszul if $\Pi_{n+1}(A)$ can be endowed with an almost Koszul grading. For an acyclic basic $n$-representation infinite algebra $A$, we show that $A$ must be Koszul if $\Pi_{n+1}(A)$ can be endowed with a Koszul grading. From this we deduce that a higher preprojective grading of an (almost) Koszul algebra $R=\Pi_{n+1}(A)$ is, in both cases, isomorphic to a cut of the (almost) Koszul grading. Up to a further assumption on the tops of the degree 0 subalgebras for the different gradings, we also show a similar result without the basic assumption in the $n$-representation infinite case. As an application, we show that $n$-APR tilting preserves the property of being Koszul for $n$-representation infinite algebras.
@@ 1. Introduction
In this way, $\Pi(Q)$ inherits a grading from tensor degrees, and one recovers $kQ$ as the degree 0 part of this grading. Furthermore, a change of orientation of $Q$ amounts to a change of the grading on $\Pi(Q)$, so while $\Pi(Q)$ as an ungraded algebra does not retain this information, the additional structure does. It is this perspective on the preprojective algebra that was generalised by Iyama and Oppermann [IO13] in Iyama’s higher Auslander-Reiten theory. Here, for a finite dimensional algebra $A$ of global dimension $n$, the $(n+1)$-preprojective algebra is defined as
@@
It is this interplay between the two gradings that we want to investigate in this article, both for the $n$-representation finite and the $n$-representation infinite case. The motivating question, which is made precise in 2.24, asks whether a given higher preprojective grading and a Koszul grading on some algebra $R$ need to be compatible in the way described above. We begin by explaining that this is not literally true and give some easy examples. However, the examples suggest that for two “incompatible” gradings we can find an automorphism that moves the higher preprojective grading to one that is compatible with the given Koszul grading. We give a proof of this statement in the $n$-representation finite case in 4.3. In the $n$-representation infinite case, we prove an analogous statement for basic acyclic algebras in 5.16, and a non-basic version for acyclic algebras over a fixed semisimple base ring in 5.17. The proof is based on the fact that if $A$ is ring-indecomposable and $n$-hereditary, and $\Pi_{n+1}(A)$ can be endowed with some (almost) Koszul grading, then $A$ can be endowed with a Koszul grading. In this way, our results can be seen as a converse to Grant-Iyama’s work, showing that $A$ can be given a Koszul grading if and only if $\Pi_{n+1}(A)$ can be given an (almost) Koszul grading. Additionally, this answers a question raised in [DG24a, Question 3.13] in the basic, $n$-representation tame case, which also applies to [DG24b]. In the $n$-representation tame case, our results have a clear geometric interpretation. The compatibility of gradings becomes equivalent to the commutation of two $k^{\ast}$-actions on the affine variety $\operatorname{Spec}(Z(\Pi_{n+1}(A)))$, as outlined in Remark 2.25. Our results also complement the work in [Thi20], where Thibault shows that certain Koszul algebras can not arise as higher preprojective algebras. Similarly, this shows that in order to construct preprojective gradings on Koszul quiver algebras, one can exhaust all gradings up to isomorphism by grading the quiver, as has been done in [HIO14, DG24a, DG24b, Gio19]. As an application, we show that $n$-APR tilting preserves Koszulity.
@@ 2.2. Graded algebras
For a graded ring $R_{\bullet}$ and graded modules $M$, NN, we denote by
@@
the graded Hom- and Ext-spaces. That means, $\operatorname{Hom}_{R_{\bullet}}(M,N)$ consists of homogeneous morphisms of degree 0.
@@ 2.2.1. Koszul algebras
[MV07, Theorem 2.8] Let $R=\bigoplus_{i\geq 0}R_{i}$ be a locally finite dimensional, nonnegatively graded algebra with $R_{0}$ semisimple. Then $R$ is Koszul if and only if $\operatorname{Ext}^{*}_{R}(R_{0},R_{0})=\bigoplus_{i\geq 0}\operatorname{Ext}^{i}_{R}(R_{0},R_{0})$ is generated in cohomological degrees 0 and $1$.
@@
Let $A$ be a finite dimensional algebra, and let $S=A/\operatorname{rad}(A)$. Then $A$ admits a Koszul grading if and only if $\operatorname{Ext}^{*}_{A}(S,S)$ is generated in cohomological degrees 0 and $1$.
@@
Consider the polynomial ring $R=k[x]=k[x+1]$, with two gradings placing $x$ respectively $x+1$ in degree $1$. Then the degree 0 parts are not isomorphic as simple $R$-modules. Of course, the automorphism exchanging $x$ and $x+1$ is a graded isomorphism for the two gradings, so we can adjust one grading so that the degree 0 parts coincide.
@@ 2.4. The Grant-Iyama construction
[GI20, Proposition 4.16] If $A_{\bullet}$ is Koszul, then the $A$-bimodule $\operatorname{Ext}^{n}_{A}(D(A),A)$ is generated in degree n-n with respect to the Koszul grading.
@@ 2.5. Higher preprojective cuts
[…]
@@
[…]
@@
[…]
@@
Denote by $E=\operatorname{Ext}^{n}_{A}(D(A),A)$ the bimodule which generates the preprojective algebra. Then it follows from 2.18 that $E$ as a graded $A$-module is generated in degree n-n. One sees easily that the Koszul degree 0 piece is $R_{0}=A_{0}$, and that $R_{1}=A_{1}\oplus E_{-n}$. Since $A$ is Koszul, it follows that $A$ is generated in degree $1$, and $E$ is generated over $A$ by $R_{1}\cap E=E_{-n}$. ∎
@@
In particular, the Koszul degree $1$ piece $R_{1}=\langle a,b,a^{\ast},b^{\ast}\rangle$ differs from $\langle a,b,(a^{\ast})^{r},(b^{\ast})^{r}\rangle$, but declaring $\{a,b\}$ to have preprojective degree $1$ still defines a preprojective grading. Denote its degree 0 part by $B=\langle e+a,f-a,(a^{\ast})^{r},(b^{\ast})^{r}\rangle$. Then even the semisimple base rings differ, since $R_{0}\cap B=0$. However, we note that the graded radical for the Koszul grading is $\operatorname{grad}(R_{\bullet})=(a,b,a^{\ast},b^{\ast})$. This equals the graded radical of the newly defined higher preprojective grading, which is given by the ideal $(a,b,(a^{\ast})^{r},(b^{\ast})^{r})$. The equality is not a coincidence, as 5.6 shows.
@@
It was proven in [HIO14] that a ring-indecomposable $n$-hereditary algebra is either $n$-representation finite or $n$-representation infinite. This is reflected in the dimension of the higher preprojective algebra being finite or infinite. We therefore split the above question into the two cases of finite and infinite as well. We will give a positive answer to this question for the $n$-representation finite case. We also prove this for basic algebras in the $n$-representation infinite case. The proof hinges on the fact that the preprojective degree 0 part can be given a Koszul grading, which may differ from the (almost) Koszul grading on the preprojective algebra.
@@ 3. Reduction to the graded basic case
[…]
Note that $e$ is homogeneous of degree 0. Therefore, $eRe$ is graded, and $eR$ is a graded $eRe$-module. This way, $\operatorname{End}_{eRe}(eR)$ becomes graded and $R_{\bullet}\simeq\operatorname{End}_{eRe}(eR)$ is a graded isomorphism. Next, we decompose the $eRe$-module eReR further as $eR=\bigoplus_{i,j}eRe_{i,j}$ according to the decomposition of $1$ in $R_{0}$. From the isotypic decomposition, we obtain that there exist elements $E_{i,j_{1},j_{2}}\in R_{0}$ such that $e_{i,j_{1}}E_{i,j_{1},j_{2}}=e_{i,j_{2}}$. Since $E_{i,j_{1},j_{2}}\in R_{0}$, we conclude that the $eRe$-modules $eRe_{i,j_{1}}$ and $eRe_{i,j_{2}}$ are graded isomorphic. Hence, we can collect the graded modules $\bigoplus_{j=1}^{d_{i}}eRe_{i,j}\simeq(eRe_{i,1})^{d_{i}}$ and obtain the graded isomorphism
@@
Let $e\in R_{0}$ be the Morita idempotent that defines the graded basic algebra $R_{\bullet}^{b}=eRe$. Decompose $e=e_{1}+\ldots+e_{m}$ inside of $R_{0}$ into primitive idempotents. It then follows from 3.6 that we have a graded isomorphism $R\simeq\operatorname{End}_{eRe}(\bigoplus_{i=1}^{m}(eRe_{i})^{d_{i}})$ for some multiplicities $d_{i}$. Since $eRe$ is graded basic, we have that $eR_{0}e$ is basic and finite dimensional. The second component of the bigrading $(\bullet,\ast)$ on $eRe$ defines a grading on $eR_{0}e$. However, it follows from 3.3 that this grading is trivial on $eR_{0}e$, and hence we have $R^{b}_{(0,0)}=\bigoplus_{i\geq 0}R^{b}_{(0,i)}$. Recall that the identity in $eRe$ is $e$, which consequently has bidegree $(0,0)$, as do all the $e_{i}$. It follows that each projective $R^{b}_{\bullet}$ module $eRe_{i}$ is bigraded. Hence, $R\simeq\operatorname{End}_{eRe}(\bigoplus_{i=1}^{m}(eRe_{i})^{d_{i}})$ becomes $\mathbb{Z}^{2}$-graded as well. Finally, it follows from 3.6 that the first component of the bigrading on $\operatorname{End}_{eRe}(\bigoplus(eRe_{i})^{d_{i}})$ recovers the $\bullet$-grading on $R$. ∎
@@
The fact that the second component of the grading $R_{\bullet,\ast}$ is a higher preprojective grading follows from [Thi20, Proposition 4.12], whose proof applies for both the $n$-representation finite and the $n$-representation infinite case. To see that $R_{\ast}$ is a cut for the Koszul grading $R_{\bullet}$, we first note that clearly $R_{1}=\bigoplus_{j\geq 0}R_{(1,j)}$ decomposes into preprojectively homogeneous pieces. Furthermore, recall that any idempotent $e\in R_{0}$ affording the Morita equivalence with $R_{\bullet}^{b}$ is full. Therefore, it follows that $R_{(1,j)}=0$ for all $j\geq 2$, because the same is true for the cut on $R^{b}_{\bullet}$. To check the second property of a cut, note that $R_{0}$ becomes a graded algebra with respect to the $\ast$-component of the bigrading. By 3.3, it follows that $R_{0}=R_{(0,0)}$, so we have that $R_{0}\subseteq\bigoplus_{i\geq 0}R_{(i,0)}$ has preprojective degree 0. Since $R_{\bullet}$ is Koszul, it is generated in degrees 0 and $1$, and hence we have that $\bigoplus_{i\geq 0}R_{(i,0)}$ is generated by $R_{(0,0)}\oplus R_{(1,0)}=R_{0}\oplus R_{(1,0)}$. ∎
@@
Suppose $R_{\bullet}$ and $_{\bullet}R$ are a Koszul and a higher preprojective grading respectively, so that $A={}_{0}R$ is $n$-representation infinite. We may choose $e\in R_{0}$ for the reduction to the graded basic algebra $R_{\bullet}^{b}$. However, in general $e\not\in{}_{0}R$ need not be preprojectively homogeneous. This means that the subspaces $e({}_{i}R)e$ need not define a grading on $eRe$, so we can not transfer the higher preprojective grading this way. In order to remedy this, we set $f={}_{0}e$ and in this way obtain a full, homogeneous idempotent in $_{0}R$, which gives rise to a graded basic algebra $fRf=\bigoplus_{i\geq 0}(f({}_{i}R)f)$. However, since $R$ is infinite dimensional, the reductions $eRe$ and $fRf$ need not be isomorphic. We therefore will assume that $R_{0}\subseteq{}_{0}R$ in 5.17, so that the idempotent $e\in R_{0}$ is also homogeneous with respect to the preprojective grading. As pointed out in Example 2.23, this assumption is not fulfilled in general. In the spirit of the Wedderburn-Malcev theorem, one can hope that $R_{0}$ can be moved into $_{0}R$ by some automorphism, so we record the following question.